OpenStudy (llort):

Where is the logical flaw? a=b a+a=a+b 2a=a+b 2a-2b=a-b-2b 2(a-b)=a-b 2=1

OpenStudy (anonymous):

fourth line, I think

OpenStudy (zarkon):

a-b=0 You can't divide by zero

OpenStudy (anonymous):

in the last step you divide both sides by (a-b) but because a=b, a-b=0 and you can't divide by 0

OpenStudy (anonymous):

in fourth line isn't that: 0 = negative number

OpenStudy (zarkon):

there looks to be a typo in line 4 also you should have 2a-2b=a+b-2b which gives you 2a-2b=a-b

OpenStudy (anonymous):

I think that's where the logical flaw is

OpenStudy (zarkon):

With that typo there are two logical flaws

OpenStudy (anonymous):

4th line a+b becomes a-b

OpenStudy (anonymous):

then division by zero